### v1

# Optimised neutron yield calculations from ($\alpha,n$) reactions with the modified SOURCES4 code

Abstract
The sensitivity of underground experiments searching for rare events such as dark matter, neutrino interactions or several beyond the standard model phenomena is often limited by the background caused by neutrons from spontaneous fission and ($\alpha,n$) reactions. A number of codes exist to calculate neutron yields and energy spectra due to these processes. In this paper we present new calculations of neutron production using the modified SOURCES4A code with recently updated cross-sections for ($\alpha,n$) reactions and the comparison of the results with available experimental data. The cross-sections for ($\alpha,n$) reactions in SOURCES4 have been taken from reliable experimental data where possible, complemented by the calculations with EMPIRE 2.19/3.2.3, TALYS 1.96 or evaluated data library JENDL-5 where the data were scarce or unavailable.
keywords:
Radioactivity , Neutron production , ($\alpha,n$) reactions , Underground experiments , Neutron background
1 Introduction
[…]
There are several computer codes to calculate the neutron yields and energy spectra of these processes [1, 2, 3, 4, 5]. Spontaneous fission is well described by the parameterisation suggested in Ref. [6] with parameters tuned to the measurements and does not depend on the material where the neutron emission occurs, but only on the concentration of 238U since other radioactive isotopes give negligible neutron yield from this process. The main problem is then the accurate calculation of ($\alpha,n$) reactions where alpha particles are produced in the decay chains of radioactive isotopes of uranium and thorium which can be found as impurities inside active media of detectors, in other detector components or in the environment. The codes use cross sections of these reactions and transition probabilities to excited states as input together with energy losses of alphas as they travel through the material until they stop.
A comparison between different codes to calculate neutron yields and spectra and experimental data has been included in several papers, see for instance Refs. [3, 4, 1, 7, 8, 9]. In this paper, we report on the new calculations of neutron production with the SOURCES4 code that includes ‘optimised’ cross-sections and transition probabilities to excited states. This ‘optimisation’ includes a combination of recent experimental data for the cross-sections with a model where the data are not available. We focus here on ($\alpha,n$) reactions caused by alphas from the radioactive decay chains of 235U, 238U, and 232Th as the main contributors to the neutron background in low-background experiments.
In the improved SOURCES4 code, we have added new experimental cross-sections, as well as cross-sections calculated with the most recent version of the TALYS code and those from the JENDL library, new excited levels and more accurate branching ratios. This paper discusses the optimisation of cross-sections implemented in SOURCES4 and compares the calculated neutron yields with experimental data from alpha beams and the entire decay chain of U and Th. Additionally, a few other modifications to the SOURCES4 and TALYS codes are discussed. With the newly implemented and optimised values, we have obtained more accurate neutron yields for most of the isotopes and compounds of interest in dark matter and neutrino experiments.
In Section 2 we briefly describe the SOURCES4 code and the cross-sections used. Section 3 includes the discussion of different sets of cross-sections for several isotopes. In Section 4 the comparison of the SOURCES4 output with experimental data is presented for some materials. Conclusions are given in Section 5.
2 The SOURCES4 code to calculate neutron production in ($\alpha,n$) reactions
The nuclear physics code SOURCES4 [1] has been used for a long time in a number of applications. It has several advantages over other similar techniques, such as flexible libraries of cross-sections and branching ratios, allowing for easy customization and updates. The code is capable of very fast calculations, providing total neutron spectra, spectra from interactions on individual isotopes, and spectra from various radioisotopes in a single calculation. It can also offer spectra from both the ground state and different excited states, making it versatile for different types of analyses. The code does not generate gamma rays from the de-excitation of final state nuclei and it cannot handle ’surface’ contamination or related problems, which is a common limitation among similar codes.
The most recent version of the code is SOURCES4C [1] but for historical reasons, we use the older version SOURCES4A [10]. Prior evaluations and release notes confirmed that the results obtained from both versions are the same when identical cross-sections and transition probabilities to the excited states are used in both versions. The simplicity of the code allows the user to select the cross-section and branching ratios of the ($\alpha,n$) reaction for a specific isotope in a material. Additionally, the user has the capability to expand the library by adding more cross-sections and branching ratios.
The original code calculates neutron production for alpha energies up to 6.5 MeV and is not fully suitable for the calculation of neutrons from radioactive processes that involve alphas with energies up to about 9 MeV. The original code SOURCES4A was modified to remove the 6.5 MeV energy cut and an updated version now allows calculations of neutron production from alphas with energies up to about 10 MeV [11, 12]. Libraries of cross-sections and branching ratios were updated with calculated values from the TALYS 1.96 [13], and EMPIRE 2.19 and 3.2.3 [14] nuclear reaction codes, as well as JENDL-5 database [15], and extended to alpha energies up to 10 MeV (see [11, 12, 16, 17, 18] for early updates). Cross-sections and branching ratios were calculated with an 0.1 MeV step in alpha energy. All these changes have been made to the version SOURCES4A and, since no changes in results were observed with the newer version of the code, SOURCES4A is continued to be used in most calculations. A comparison of cross-sections from EMPIRE 2.19 with some experimental data was published in Refs. [12, 18] and the results from the modified SOURCES4A code were used in a number of dark matter experiments (see, for example, Refs. [19, 20, 21, 22]). We have recently modified the code a) to allow calculation of neutron yields and spectra from different materials in one run; b) to allow the choice of the cross-section and branching ratios in the input file without the need to change positions of the cross-section or transition probabilities in the libraries (tape 3 and tape 4); c) to increase the maximum number of target elements to 100; d) to increase the maximum number of discrete nuclear levels to 500.
The user input for SOURCES4A includes either the energy of an alpha particle or the $Z$ and $A$ of the radioactive isotope (or several isotopes in the case of decay chains, for instance) with the number of radioactive atoms per unit volume. The user should also specify material composition (where the alpha sources are located) and isotopic composition for each element (only isotopes with cross-sections present in the code library can be included).
A comparison of cross-sections from TALYS 1.9, EMPIRE 2.19/3.2.3, and experimental data was presented in Refs. [23, 24]. These papers also included some preliminary results on the comparison of neutron yields calculated using SOURCES4A with different cross-sections taken from TALYS 1.9, EMPIRE 2.19/3.2.3, and some experimental data.
In this work we report the very latest amendments to the libraries of SOURCES4A that include cross-sections calculated with the latest version of TALYS 1.96 [13] and evaluated data from JENDL-5 [15], as well as some very recent experimental data. We have also optimised the choice of cross-sections for neutron yield calculations to get a better agreement with experimental data including those from alpha beams.
The TALYS and EMPIRE nuclear reaction codes based on statistical models provide an accurate description of reaction mechanisms over wide energy and nuclear mass ranges. These codes return reliable results for ($\alpha,n$) reactions within the energy range up to 200 MeV, for atomic numbers greater than 9 and mass numbers between 20 and 339. More precisely, the codes are reasonably accurate for isotopes which do not have resonances in the energy dependent cross-sections. For light elements with a large number of resonances experimental data or evaluated data from the JENDL-5 library are to be used to achieve better accuracy. The alpha-particle sub-library in JENDL-5 provides evaluated cross-section data for alpha-induced reactions on 18 light nuclides from 6Li to 30Si, with an upper limit of 15 MeV for the incident energy.
In the default version of the TALYS 1.96 nuclear reaction code, the maximum number of discrete levels that are considered in the Hauser-Feshbach model for residual nuclei is set to 30. We have modified the TALYS code in order to increase the number of levels to include all the possible excited states for the final-state nucleus that are available in RIPL-3 [25] library.
Section 3 presents a comparison of our ’optimised’ cross-sections with measurements and results from the statistical codes TALYS and EMPIRE, along with data from the JENDL database, as a few examples relevant to underground experiments aimed at detecting rare events.
3 Cross sections comparison
The cross-sections from different models and experimental data were compared in Ref. [23]. The same paper included calculations of neutron yields with SOURCES4A using different models for the cross-sections. Here we extend previous work and present neutron yield calculations with optimised cross-sections as explained in Section 2 comparing this to available experimental data. We focus on materials extensively used in rare event physics experiments, such as those searching for dark matter, neutrino interactions, or other phenomena beyond the Standard Model.
The list of other elements considered in this work and included in the library of SOURCES4A comprises Be, B, N, Na, Ca, P, Cl, Mg, Ti, Fe, Ni, Cr, Mn, Cu and others.
3.1 Lithium
Recently, the CRESST Collaboration suggested the idea of using LiAlO2 crystals for direct dark matter searches due to the fact that 7Li is a good candidate for studying spin-dependent dark matter interactions in the sub-GeV mass region. 6Li also has a high cross-section for neutron capture providing a unique signature that allows monitoring of the neutron flux on site [26, 27].
Excitation functions for 6Li($\alpha,n$)9B and 7Li($\alpha,n$)10B reactions from threshold to 15 MeV were reported in Ref. [28] and included in JENDL-5 library. Total cross-sections were evaluated to match the measurements of thick target neutron yields from Ref. [29]. These cross-sections are shown in Figure 1(b)) together with other data available for 7Li.
Figure 1: (a) 6Li($\alpha$,n)9B cross section as a function of alpha energy. (b) 7Li($\alpha$,n)10B cross section as a function of alpha energy. Experimental data are taken from [29]. Uncertainties are not shown on this and other figures to make data points and calculations clearly visible; these uncertainties can be found in the original publications.
3.2 Carbon
Carbon contains 1.07% of 13C and is present in scintillators and various plastics and other polymers, such as acrylic, polyethylene, nylon and PTFE. These materials are often placed near the target volume in rare event experiments. In addition, carbon is prevalent in the composition of rocks.
The most recent experimental data available for the cross sections from Ref. [30] are only for the ground state population. Thus, for this isotope, the measurements from Ref. [31] are used in this work up to $\sim 8$ MeV and complemented with TALYS 1.96 calculations above this energy (see Figure 2).
Figure 2: 13C($\alpha$,n)16O cross section as a function of alpha energy. Experimental data are taken from [30, 31, 32, 33, 34, 35, 36, 37]. 12C does not contribute to neutron yield since the energy threshold of the corresponding reaction exceeds the maximum energy of alphas from radioactive decay chains of natural uranium and thorium.
3.3 Oxygen
Oxygen is present in water, water-based liquid scintillators (WbLS) [38], some plastics, and rocks. Two isotopes are of interest for neutron production in ($\alpha,n$) reactions: 17O and 18O with abundances of 0.038% and 0.205%, respectively. 16O has an energy threshold for ($\alpha,n$) reaction above the energies of alphas from radioactivity.
For 17O, the measurements from Ref. [39, 34] are used up to 5.2 MeV and complemented with the JENDL-5 data above this energy (see Figure 3(a)). In the case of 18O, the cross sections evaluated using JENDL-5 were used (see Figure 3(b)).
Figure 3: (a) 17O($\alpha$,n)20Ne cross section as a function of alpha energy. (b) 18O($\alpha$,n)21Ne cross section as a function of alpha energy. Experimental data are taken from [40, 34, 39].
3.4 Fluorine
Polytetrafluoroethylene (PTFE) is widely used in rare event experiments using xenon or argon as active medium because it withstands temperatures of liquid xenon (LXe) and liquid argon (LAr) and is a good insulator [41, 42, 43]. It is also an excellent reflector for VUV light at the LXe and LAr scintillation wavelengths. In this case, the measurements from Ref. [44] were used up to 7 MeV and complemented with TALYS 1.9 calculations above this energy. Figure 4 shows a comparison of the experimental cross-sections with calculations.
Figure 4: (a) 19F($\alpha$,n)22Na cross section as a function of alpha energy. Experimental data are taken from [44, 45, 46, 47, 48, 49].
3.5 Aluminium
Sapphire crystals containing aluminium have been proposed for light dark matter detection [50, 51]. Aluminium is also used in the readout electrodes and other components.
For aluminium, the cross-sections measured in Refs. [52, 53] were used and complemented with those calculated with EMPIRE 3.2.3 above 9.2 MeV (see Figure 5). This energy is already above the highest alpha energy from radioactive decay chains.
Figure 5: 27Al($\alpha$,n)30P cross section as a function of alpha energy. Experimental data are taken from [54, 53, 52].
3.6 Silicon
Silicon crystals of high purity are employed as substrates for detectors in dark matter direct detection experiments, such as CDMS [55], CDMS-II [56] and SuperCDMS [57]. Silicon is also present in quartz, glass and light sensors.
For 29Si, the measured cross-section from Ref. [52] was used up to 6.8 MeV and replaced with the cross-section from TALYS 1.96 above this energy (see Figure 6). As discussed above, none of the statistical models used in the codes can accurately predict the resonance shape of the cross-section and the use of experimental data is unavoidable. Experimental data sets meanwhile are consistent and can be used up to energies where they are available (6.8 MeV). Above this energy, the choice of the cross-section is quite arbitrary and an option of using TALYS 1.96 was dictated by: a) the most recent update to the calculations and b) a reasonable agreement with experimental data at energies close to the upper end of data range.
Figure 6: 29Si($\alpha$,n)32S cross section as a function of alpha energy. Experimental data are taken from Refs. [52, 58, 59].
3.7 Noble elements used in neutrino and dark matter experiments
Among the noble elements, argon and xenon are the only ones currently used as targets for neutrino and dark matter experiments due to their high scintillation yield, low intrinsic radioactivity, and excellent electron transport characteristics [60]. Neon has been proposed for detecting low-energy neutrinos and might also be used for WIMPs searches [61].
For all isotopes of xenon, the energy threshold for ($\alpha,n$) reactions, which is determined by the Coulomb barrier for high-$Z$ materials, is above 9 MeV. These energies are higher than those of $\alpha$-particles from radioactive decay chains.
Argon is widely used as active volume in rare event searches and several examples include DUNE [62], DarkSide [63] and ArDM [64] For $\alpha$-particle energies up to 10 MeV, the energy dependence of the ($\alpha,n$) reaction cross-sections calculated with TALYS and EMPIRE nuclear reaction codes agree reasonably well. However, there is a significant discrepancy between the calculated values and the experimental data point. This data point has been measured at a fixed neutron scattering angle and extrapolated to the total cross-section giving potentially quite a high uncertainty of the procedure.
Figure 7: 40Ar($\alpha$,n)43Ca cross-section as a function of alpha energy. Experimental data is taken from Ref. [65].
4 Comparison of SOURCES4 with experimental data
Figures 8 and 9 show the dependence of the neutron yield on the alpha energy for two materials, as examples, that are frequently used in underground experiments, aluminium and silicon oxide. Following a similar approach as in previous publications [7, 4], the neutron yield is given as the number of neutrons per $10^{6}$ alphas. In our calculations with SOURCES4A we use optimised cross-sections - a combination of experimentally measured cross-sections and calculations or evaluations for energies where data are not available, as described in Section 3. A good agreement with measurements is observed for all materials tested confirming the right choice of cross-sections for the isotopes in these materials.
Figure 8: 27Al($\alpha$,n)30P neutron yield as a function of alpha energy. Optimised cross-sections are used in SOURCES4A. Experimental data are taken from [66, 67, 68, 69, 29].
Figure 9: SiO2 neutron yield as a function of alpha energy. Optimised cross-sections are used in SOURCES4A. Experimental data are taken from [68].
Neutron energy spectra as calculated with SOURCES4 and measured in Ref. [68] for 5 MeV $\alpha$ particles incident on fluorine and aluminium targets are shown in Figure 10(a). The shape of the spectra are reproduced correctly in the calculations and, given the difficulty to accurately reconstruct neutron energy in an experiment the overall agreement is reasonably good.
Figure 10: Neutron energy spectrum from 5 MeV $\alpha$ particles in (a) 19F and (b) 27Al in comparison with data from Ref. [68].
In underground experiments the prime interest is in the evaluation of neutron yields and energy spectra in ($\alpha,n$) reactions from naturally occurring radioactive isotopes and decay chains. A comparison of the calculations with measurements from Ref. [70] for the whole uranium and thorium decay chains is shown in Figure 11 whereas a similar comparison with evaluated data from Ref. [7] is included in Figure 12. Only experimental uncertainties are shown as quoted in Refs. [70, 7].
For most materials the agreement between calculations and data is within 10%, again confirming correct choice of cross-sections in SOURCES4 library. Neutron yields for carbon agree with the measurements reported in Ref. [70] (Figure 11) but are higher than those reported in Ref. [7]. We note that the data in Ref. [7] are not the direct measurements of neutron yields from uranium and thorium decay chains but were extrapolated from neutron yields measured with $\alpha$ particle beams taking into account the $\alpha$ decay energies and energy losses in different materials leading to an additional uncertainty in the results. Nevertheless, neutron yield in carbon may indeed be overestimated given that our calculations for 13C use the recently measured cross-section which is higher than previous measurements (see Figure 2. The biggest difference between our calculations and data reported in Ref. [7] is for 238U decay chain in iron but, surprisingly, the agreement for 232Th decay chain is pretty good, pointing probably to the additional uncertainty in the reported values in Ref. [7] due to the yield extrapolation method. Experimental data for Fe have not been quoted in that paper.
Similar calculations of neutron yields and energy spectra have been performed using the NeuCBOT code which has now an option of using either JENDL-5 or TENDL libraries [2]. Neutron yields from GEANT4 and associated SaG4n code using two libraries JENDL-5 [71] and TENDL-2017 [72] were reported in Ref. [4]. A comparison between calculations with the optimised SOURCES4 and these codes for a few materials was included in Ref. [9].
Figure 11: Ratio of our calculations with optimised SOURCES4 to the neutron yield measured in Ref. [70] for: (a) natU and (b) 232Th decay chains in several materials. The shaded area represents the ±$\pm 10\%$ difference. Only the experimental errors are included.
Figure 12: Ratio of the calculated values from the optimised SOURCES4 code to the evaluated neutron yield data from Ref. [7] for (a) 238U and (b) 232Th decay chains in several materials. The shaded area represents the ±$\pm 10\%$ difference. No error was given for Fe. Only the experimental errors were taken into account for other materials.
Table 1: Neutron yield from ($\alpha,n$) reactions originated in natU and 232Th decay chains in several materials as calculated by the SOURCES4 code. Neutron yield is given as the number of neutrons per gram of material per second per ppb of U or Th concentration. Spontaneous fission of 238U gives a neutron yield of $1.35\times 10^{-11}$ n/g/s/ppb in all materials.
Element natU 232Th Compound natU 232Th
Li $7.12\times 10^{-10}$ $2.95\times 10^{-10}$ Al2O3 $8.53\times 10^{-11}$ $4.17\times 10^{-11}$
Be $8.38\times 10^{-9}$ $2.79\times 10^{-9}$ BeO $3.08\times 10^{-9}$ $1.03\times 10^{-9}$
B $1.99\times 10^{-9}$ $6.14\times 10^{-10}$ C2F4 $9.76\times 10^{-10}$ $3.90\times 10^{-10}$
C $1.76\times 10^{-11}$ $7.04\times 10^{-12}$ CaCO3 $7.28\times 10^{-12}$ $2.89\times 10^{-12}$
N $5.80\times 10^{-11}$ $3.23\times 10^{-11}$ CaF2 $6.63\times 10^{-10}$ $2.75\times 10^{-10}$
Na $4.13\times 10^{-10}$ $1.93\times 10^{-10}$ CH2 $1.71\times 10^{-11}$ $7.04\times 10^{-12}$
Mg $2.03\times 10^{-10}$ $7.67\times 10^{-11}$ H2O $3.98\times 10^{-12}$ $1.39\times 10^{-12}$
Al $1.67\times 10^{-10}$ $8.25\times 10^{-11}$ H3BO3 $3.38\times 10^{-10}$ $9.64\times 10^{-11}$
Si $2.18\times 10^{-11}$ $1.01\times 10^{-11}$ MgO $1.20\times 10^{-10}$ $4.56\times 10^{-11}$
P $2.85\times 10^{-11}$ $1.94\times 10^{-11}$ Na2CO3 $2.78\times 10^{-10}$ $1.29\times 10^{-10}$
Cl $8.08\times 10^{-11}$ $4.36\times 10^{-11}$ NaCl $1.52\times 10^{-9}$ $6.07\times 10^{-10}$
Ar $1.52\times 10^{-10}$ $9.00\times 10^{-11}$ NaF $8.16\times 10^{-10}$ $3.50\times 10^{-10}$
Ca $1.80\times 10^{-12}$ $1.22\times 10^{-12}$ PbF2 $4.39\times 10^{-10}$ $1.74\times 10^{-10}$
Ti $3.63\times 10^{-11}$ $3.18\times 10^{-11}$ SiO2 $1.41\times 10^{-11}$ $5.98\times 10^{-12}$
Cr $1.40\times 10^{-11}$ $1.45\times 10^{-11}$ Stainless steel Fe(66%), Cr(17%), Ni(12%), Mn(2%), Mo(2%), Si(1%) $7.38\times 10^{-12}$ $8.91\times 10^{-12}$
Mn $9.29\times 10^{-12}$ $1.04\times 10^{-11}$
Fe $4.74\times 10^{-12}$ $6.68\times 10^{-12}$
Ni $1.02\times 10^{-13}$ $2.63\times 10^{-13}$ UC $2.50\times 10^{-12}$ $1.02\times 10^{-12}$
Cu $3.67\times 10^{-13}$ $1.07\times 10^{-12}$ UO2 $2.41\times 10^{-12}$ $8.37\times 10^{-13}$
5 Conclusions
The optimised approach for calculating neutron yields using SOURCES4 code shows a good agreement with experimental data, proving the importance of using reliable cross-section data where available. In cases where experimental data are not available for a specific isotope, nuclear reaction codes such as TALYS, EMPIRE, or databases like JENDL provide a good alternative. The choice of models for the cross-sections is confirmed by comparing calculated neutron yields with alpha beam data to ensure accuracy and reliability. This approach ensures that the most accurate and recent information is used for neutron production calculations, which is very important for applications in rare event physics experiments. By increasing the number of discrete nuclear levels considered in TALYS 1.96, the modified code now accounts for all possible excited states for the final state nucleus, as available in RIPL-3.
6 Acknowledgements
V.K. and P.K. would like to thank the Science and Technology Facilities Council (STFC, UK) and the University of Sheffield for financial support. For M.P. this work was performed with the financial support of the Romanian Program PNCDI III, Programme 5, Module 5.2 CERN-RO, under contract no. 04/2022.